OpenStudy (anonymous):

Divide (10 × 1017) ÷ (100 × 1013). Express the answer in scientific notation

OpenStudy (diyadiya):

\[\Large \frac{10 \times 10^{17}}{100 \times 10^{13}}\] Remember this \[\Large x^m \times x^n = x^{m+n}\]\[\Large \frac{x^m}{ x^n} = x^{m-n}\]

OpenStudy (diyadiya):

No to make it easier ; do you know what is \(\large 10^2 \) ??

OpenStudy (anonymous):

100

OpenStudy (diyadiya):

Ok now the denominator can be written as \(\large 10^2 \times 10^{13} \) ok?

OpenStudy (diyadiya):

did you understand this part?

OpenStudy (anonymous):

yea but i cant get what 10^13 is

OpenStudy (diyadiya):

Its there in your Question!

OpenStudy (diyadiya):

\[\Large \frac{10 \times 10^{17}}{100 \times 10^{13}} = \Large \frac{10 \times 10^{17}}{10^2 \times 10^{13}}\]

OpenStudy (diyadiya):

Is this part clear?

OpenStudy (anonymous):

1.00 × 10^3 1.00 × 10^30 0.10 × 10^4 0.10 × 10^30 these are the answers i can pick from

OpenStudy (diyadiya):

well, i can't pick an answer for you but i can help you to get one ;)

OpenStudy (anonymous):

i know

OpenStudy (diyadiya):

What is 10 x \(10^{17} \) = ?

OpenStudy (diyadiya):

\[\large 10 \times 10^{17} = 10^1 \times 10^{17}= 10^{1+17}= 10^{18}\] i used this formula \[\Large x^m \times x^n = x^{m+n}\]

OpenStudy (diyadiya):

Now can you tell me what is \(\large 10^2 \times 10^{13} \)=?

OpenStudy (anonymous):

10^15

OpenStudy (diyadiya):

Right! now we have \[\Large \frac{10^{18}}{10^{15}}\] Now use this formula \[\Large \frac{x^m}{ x^n} = x^{m-n}\] what will you get ?

OpenStudy (anonymous):

1^3

OpenStudy (diyadiya):

1^3 ?

OpenStudy (diyadiya):

you mean 10^3?

OpenStudy (anonymous):

so u dont do 10 divided by 10

OpenStudy (diyadiya):

Nope as i said \[\Large \frac{x^m}{ x^n} = x^{m-n} \] For example :\[\Large \frac{10^5}{10^2}= 10^{5-2} = 10^3\] Just like we did before \[\Large 10^5 \times 10^2 = 10^{5+2}= 10^7\] When we have to multiply with same bases we just add the powers & when its division we subtract the powers so \[\large \frac{10^{18}}{10^{15}} = 10^{18-15}=10^3\] got this much?

OpenStudy (anonymous):

got it

OpenStudy (diyadiya):

ok so according to your choices your answer will be 1.00 x \(10^3\) since 1.00 is same as 1 and when you multiply 1 with any number you get the same number. understood?

OpenStudy (anonymous):

thank u so much

OpenStudy (diyadiya):

Anytime :)

OpenStudy (diyadiya):

Remember both the formulas ,It will be usefull for similar problems (There's nothing to memorise anyways, its easy :D )

OpenStudy (anonymous):

i am doing an online class and there is nobody to show me how it is done. and no 2 problems are alike

OpenStudy (diyadiya):

You can use Openstudy! to clear your doubts :)

OpenStudy (anonymous):

do u have time to help me with one more

OpenStudy (anonymous):

Multiply (3.6 × 10^15) × (8 × 10^21). Express the answer in scientific notation

OpenStudy (diyadiya):

First do 3.6 x 8

OpenStudy (anonymous):

28.8

OpenStudy (diyadiya):

Now do 10^15 x 10^21

OpenStudy (anonymous):

10^36

OpenStudy (diyadiya):

Very Good! so the answer is?

OpenStudy (anonymous):

28.8X10^36

OpenStudy (diyadiya):

Well in the Scientific form we can only have 1 digit before the decimal here we have 28.8 that is 2 numbers before decimal so we multiply and divide by 10 \[\Large 28.8 \times 10^{36} \times \frac{10}{10}= \frac{28.8}{10} \times 10^{36} \times 10 = 2.88 \times 10^{36} \times 10 \] 10^36 x 10 = 10^37 so answer is 2.88 x 10^37

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which number is greater than 3.25 ×10^30. now i have some of these

OpenStudy (diyadiya):

Please post it as New question! some one else might help :) I've to leave now! Bye!